How many grams of sodium chloride (NaCl) are needed to make 345 mL of of a 2 M solution? 


1
Expert's answer
2021-04-08T12:10:34-0400

Molarity = 2M , volume = 345 mL = 0.345L

Molarity = moles of solute/volume of solution

2 = moles / 0.345

moles = 2 × 0.345 = 0.69 mol

Moles of NaCl = 0.69 mol

Gram of NaCl = 0.69×58.44 =40.32 g